OpenStudy (anonymous):

In the xy-plane, line t has equation y=1/4x+3, and line m has equation y=1/2x+b, If the lines intersect at the point with coordinates (t,11/2), what is the value of b.

OpenStudy (anonymous):

substitute y=11/2 into the equation

OpenStudy (adi3):

@ganeshie8

OpenStudy (anonymous):

\[y=\frac{ x }{ 4 }+3\]

OpenStudy (anonymous):

\[\frac{ 11 }{ 2 }=\frac{ x }{ 4 }+3\]

OpenStudy (adi3):

@AaronAndyson

OpenStudy (anonymous):

Multiply 2 on both sides

OpenStudy (anonymous):

so you're basically solving for x after that?

OpenStudy (anonymous):

\[2\times \frac{ 11 }{ 2 }=2(\frac{ x }{ 4 }+3)\]

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

sorry continue

OpenStudy (anonymous):

\[11=\frac{ x }{ 2 }+6\]

OpenStudy (anonymous):

subtract 6 on both sides

OpenStudy (anonymous):

\[5=\frac{ x }{ 2 }\]

OpenStudy (anonymous):

Then,multipy 2 on both sides

OpenStudy (anonymous):

\[x=5\times2\]\[x=10\]

OpenStudy (anonymous):

substitute x=10 and y=11/2 into the second equation in oder to find the value of b

OpenStudy (anonymous):

\[y=\frac{ x }{ 2 }+b\]

OpenStudy (anonymous):

\[\frac{ 11 }{ 2 }=\frac{ 10 }{ 2 }+b\]

OpenStudy (anonymous):

\[\frac{ 11 }{ 2 }=5+b\]

OpenStudy (anonymous):

Multiply 2 on both sides

OpenStudy (anonymous):

\[2\times \frac{ 11 }{ 2 }=2(5+b)\]

OpenStudy (anonymous):

\[11=10+2b\]

OpenStudy (anonymous):

subtract 10 on both sides

OpenStudy (anonymous):

\[11-10=2b\]

OpenStudy (anonymous):

\[1=2b\]

OpenStudy (anonymous):

divide 2 on both sides

OpenStudy (anonymous):

\[b=?\]solve it @Mhm120